# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Acknowledgements.
We acknowledge insightful discussions with, and feedback from, Adam Shaw, Ran Finkelstein, Pascal Scholl, Joonhee Choi, and Soonwon Choi. We acknowledge support from the Gordon and Betty Moore Foundation (Grant GBMF11562), the Weston Havens Foundation, the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-1733907), the NSF QLCI program (2016245), the NSF CAREER award (1753386), the Army Research Office MURI program (W911NF2010136), the U.S. Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), and the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. HJM acknowledges support from the NSF Graduate Research Fellowship Program under Grant No. 2139433. KHL acknowledges support from the AWS-Quantum postdoctoral fellowship and the NUS Development Grant AY2023/2024.

### v3

@@ Acknowledgements.
We acknowledge insightful discussions with, and feedback from, Adam Shaw, Harry Levine, Zunqi Li, Ran Finkelstein, Pascal Scholl, Joonhee Choi, Dolev Bluvstein, and Soonwon Choi. We acknowledge support from the Gordon and Betty Moore Foundation (Grant GBMF11562), the Weston Havens Foundation, the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-2317110), the NSF QLCI program (2016245), the NSF CAREER award (1753386), the Army Research Office MURI program (W911NF2010136), the U.S. Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), and the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. H.J.M. acknowledges support from the NSF Graduate Research Fellowship Program under Grant No. 2139433. K.H.L. acknowledges support from the AWS-Quantum postdoctoral fellowship and the NUS Development Grant AY2023/2024.
@@
Author contributions
H.J.M. and M.E. conceived the idea and experiment. H.J.M., G.N., E.B., K.H.L., and X.L. performed numerical simulations and contributed to the experimental setup. H.J.M., G.N., E.B., and K.H.L. performed the experiments. H.J.M., G.N., and E.B. analyzed the data. All authors contributed to the manuscript. M.E. supervised this project.
Competing interests
The authors declare no competing interests.